a student took 18 minutes to walk from home to school.given that his average speed is 4km/h,how far he had to travel
Answer by nerdybill(7384)

a student took 18 minutes to walk from home to school.given that his average speed is 4km/h,how far he had to travel
.
Apply the distance formula:
d = rt
.
However, begin by changing the 18 minutes into hours:
18min * 1hr/60min
= 18/60
= 3/10 hour
.
distance = 3/10 * 4
distance = 3/5 * 2
distance = 6/5 km
or
distance is 1.2 km